Find two numbers whose sum is 50 and difference is 12
Answer by mangopeeler07(462)

Number 1=x
Number 2=y
x+y=50
x-y=12
Eliminate y by adding the equations to each other
2x=62

Then divide both sides by 2
x=31

Plug x into the first equation.
31 +y=50.

Subtract 31 from both sides
y=19

Therefore,
x=31
y=19